# Nathalie Gimenes PUC-Rio Tonghui Qi

### v1

@@ C.3 Proofs of results
The uniform convergence is a consequence of Lemma C.3.
@@
due to the triangle inequality and bounding by an envelope of $\mathcal{Q}$. It will be instructive here to prove that $q$ is (pointwise) Lipschitz continuous with respect to $\psi$ for any $\theta$ in the following sense:
@@
for some function $c$ such that $E\left[c^{2}\right]<\infty$. The main calculation involves showing that $T\left(\psi\right)=\left(\psi_{1}\circ\psi_{2}\right)\left(1-\psi_{2}\right)$ is Lipschitz continuous, which will also be useful in other parts of the appendix, so we provide some details here.
@@
where $\left\|\epsilon\omega_{1}^{\left(1\right)}\times\epsilon\omega_{2}\right\|_{\infty}$ is the leading term, as it can be shown that $\left\|\widehat{\psi}_{1}^{\left(1\right)}-\psi_{10}^{\left(1\right)}\right\|_{\infty}=O_{p}\left(\sqrt{\frac{\log L}{Lh^{3}}}+h^{s-1}\right)$ and $\left\|\widehat{\psi}_{2}-\psi_{20}\right\|_{\infty}=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right)$; these rates correspond to the convergence rates in Proposition 4, with $j=2$, and Proposition 5 respectively. We then have $\Xi\left(\epsilon\omega\right)=O_{p}\left(\frac{\log L}{Lh^{3/2}}+h^{2s}\right)=o_{p}\left(L^{-1/2}\right)$, where the latter holds due to our assumption on $h$. This ensures that $T\left(\widehat{\psi}\right)-T\left(\psi_{0}\right)=\dot{T}\left(\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]+o_{p}\left(L^{-1/2}\right)$.
@@
Next, we re-write the integrand above as a functional of the observed bid’s quantile and its derivative,B\ B and $B^{\left(1\right)}$ respectively. Recalling the relations between bid and valuation quantiles, see (10) and (11), let us define
@@
In the proof of Lemma 5, using (26), we have shown in (27) that $q$ is Lipschitz in $\psi$ for fixed $\theta$. It is easy to extend this over $\Theta$, as $U_{\theta}^{\left(j\right)}$ is Lipschitz due to continuity of $U_{\theta}^{\left(j+1\right)}$ on $\Theta$ for $j=0,1$. It is easy to verify that,

### v2

@@ C.3 Proofs of results
The uniform convergence of $Q_{L}$ to $Q$ in probability is a consequence of Lemma C.3.
@@
due to the triangle inequality and bounding by an envelope of $\mathcal{Q}$. It will be instructive here to show that $q$ is (pointwise) Lipschitz continuous with respect to $\psi$ for any $\theta$ in the following sense:
@@
for some function $c$ such that $E\left[c^{2}\right]<\infty$. The main calculation for establishing Lipschitz continuity involves showing that $T\left(\psi\right)=\left(\psi_{1}\circ\psi_{2}\right)\left(1-\psi_{2}\right)$ is Lipschitz continuous, which will also be useful in other parts of the appendix, so we provide some details here.
@@
where $\left\|\epsilon\omega_{1}^{\left(1\right)}\times\epsilon\omega_{2}\right\|_{\infty}$ is the leading term, as it can be shown that $\left\|\widehat{\psi}_{1}^{\left(1\right)}-\psi_{10}^{\left(1\right)}\right\|_{\infty}=O_{p}\left(\sqrt{\frac{\log L}{Lh^{3}}}+h^{s-1}\right)$ and $\left\|\widehat{\psi}_{2}-\psi_{20}\right\|_{\infty}=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right)$. These rates correspond to the convergence rates in Proposition 4 with $j=2$ and Proposition 5 respectively. We then have $\Xi\left(\epsilon\omega\right)=O_{p}\left(\frac{\log L}{Lh^{3/2}}+h^{2s}\right)=o_{p}\left(L^{-1/2}\right)$, where the latter holds due to our assumption on $h$. This ensures that $T\left(\widehat{\psi}\right)-T\left(\psi_{0}\right)=\dot{T}\left(\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]+o_{p}\left(L^{-1/2}\right)$.
@@
Next, we re-write the integrand above as a functional of the observed bid’s quantile and its derivative,B\ B and $B^{\left(1\right)}$ respectively. Let us define
@@
In the proof of Lemma 5, using (25), we have shown in (26) that $q$ is Lipschitz in $\psi$ for fixed $\theta$. It is easy to extend this over $\Theta$, as $U_{\theta}^{\left(j\right)}$ is Lipschitz due to continuity of $U_{\theta}^{\left(j+1\right)}$ on $\Theta$ for $j=0,1$. It can then be verified that,